OpenStudy (anonymous):

A town has a clock tower with a bell that tolls the number of hours of the day. (The bell will toll six times at 6 a.m. and six times at 6 p.m.) How many times in a day will the bell toll?

OpenStudy (anonymous):

24 `sum_(i=1)^24 i` `2sum_(i=1)^12 i` 24!

OpenStudy (anonymous):

these are my options

OpenStudy (fibonaccichick666):

To start what does 24! represent?

OpenStudy (anonymous):

i have no idea all that i know these are my option for answers

OpenStudy (fibonaccichick666):

yes, 24! is one of the answers. But I am curious to see if you know this concept before we examine the summations

OpenStudy (anonymous):

never mind i figure it out but thanks anyways